# The cohomological Kudla conjecture for unitary Shimura varieties

### v1

@@ 5. Proofs of the Main Theorems
In this section, we prove Theorem 1.4 and Theorem 1.3.
Let $(L,h)$ be a Hermitian lattice of signature $(n+1,1)$ over ${\mathcal{O}}_{k}$, the ring of integers of a quadratic imaginary field $k$ and let $X^{\mathrm{tor}}_{\Gamma}$ be the toroidal compactification of the unitary Shimura variety constructed in Section 3.
Let $g\geq 1$, $N\in\mathrm{Herm}_{g}(k)_{\geq 0}$ and $\underline{\nu}\in(L^{\vee}/L)^{g}$. Let ${\mathcal{Z}}(\underline{\nu},N)$ be the associated special cycle. We still denote its closure in $X^{\mathrm{tor}}_{\Gamma}$ by $\overline{{\mathcal{Z}}}(\underline{\nu},N)$.
For each ${\mathfrak{J}}\subset L$ isotropic rank $1$ ${\mathcal{O}}_{k}$-submodule, let $M={\mathfrak{J}}^{\bot}/{\mathfrak{J}}$ and let ${\mathcal{B}}_{\mathfrak{J}}$ be the corresponding boundary divisor of $X^{\mathrm{tor}}_{\Gamma}$. By Lemma 3.5, the restriction of Kudla–Millson generating series to ${\mathcal{B}}_{\mathcal{J}}$ is the generating series of the $Z(\underline{\lambda})$ cycles introduced in Section 3.5:
$\Phi_{M}^{g}(\tau)=\sum_{\underline{\lambda}\in(M^{\vee})^{g}}[Z(\underline{\lambda})]q^{h(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\leavevmode\nobreak\ .$
We know define the completions of the special cycles: For each ${\mathfrak{J}}\subset L$ isotropic ${\mathcal{O}}_{k}$-line, let $M={\mathfrak{J}}^{\bot}/{\mathfrak{J}}$, which is a Hermitian lattice of signature $(n,0)$. For each $\underline{\nu}\in(M^{\vee}/M)^{g}$ and $N\in\mathrm{Herm}_{g}(k)_{>0}$, we define:
$P_{i}^{\ell}(\underline{\nu},N)=\sum_{\underset{[\underline{\lambda}]=\underline{\nu}}{\underline{\lambda}\in(M^{\vee})^{g}}}P_{i}^{\ell}(\lambda_{1},\ldots,\lambda_{g})\leavevmode\nobreak\ .$
@@
$[\varphi(\tau,\underline{\nu},N)]=\sum_{\underset{[\underline{\lambda}]=\underline{\nu}}{\underline{\lambda}\in(M^{\vee})^{g}}}\sum_{\ell=1}^{g}\frac{(-1)^{\ell}}{(4\pi)^{\ell}(g-\ell)!}\mathrm{Tr}\left(\Lambda^{\ell}(Y)\Lambda^{g-\ell}([f(\underline{\lambda})]\right)\wedge D^{\ell-1}\leavevmode\nobreak\ .$
@@
    the completed special in $X_{\Gamma}^{\mathrm{tor}}$: choose a set ${\mathcal{J}}$ of representative of isotropic lines of $L$ under the action of $\Gamma$ and let
    $[\widehat{{\mathcal{Z}}}(\underline{\nu},N,Y)]=[\overline{{\mathcal{Z}}}(\underline{\nu},N)]-\sum_{{\mathfrak{J}}\in{\mathcal{J}}}\frac{r_{\mathfrak{J}}}{d_{k}}\iota_{{\mathfrak{J}},*}[\varphi(\tau,\underline{\nu},N)]\leavevmode\nobreak\ .$
@@
    $[\widetilde{{\mathcal{Z}}}(\underline{\nu},N)]=[\overline{{\mathcal{Z}}}(\underline{\nu},N)]+\sum_{{\mathfrak{J}}\in{\mathcal{J}}}\frac{r_{\mathfrak{J}}}{d_{k}}\sum_{\underset{i\in I_{\ell}}{\ell=0}}^{g-1}P^{\ell}_{i}(\underline{\nu},N)\iota_{{\mathfrak{J}},*}[W^{\ell}_{i}\cup D^{g-\ell-1}]\leavevmode\nobreak\ ,$
where $-d_{k}$ is the discriminant of $k$ and $r_{\mathfrak{J}}$ is the integer appearing in Section 3.3.
Let $\alpha\in\mathrm{H}_{2g}(X_{\Gamma}^{\mathrm{tor}})$, and let’s prove that both pairings $(\widehat{\Phi}_{L}.\alpha)$ and $(\widetilde{\Phi}_{L}.\alpha)$ is a Hermitian modular form of weight $1+n$ and genus $g$.
By Lemma 3.3, every $\alpha\in\mathrm{H}_{2g}(X_{\Gamma}^{\mathrm{tor}})$ can be written as a sum $\alpha=\beta+\gamma$ where $\beta\in H_{2g}(X_{\Gamma},\mathbb{Q})$ and $\gamma\in\mathrm{H}_{2g}(\partial(X_{\Gamma}^{\mathrm{tor}}),\mathbb{Q})$. In particular, we have:
$\left(\widetilde{\Phi}.\alpha\right)=\left(\widetilde{\Phi}.\beta\right)+\left(\widetilde{\Phi}.\gamma\right)\leavevmode\nobreak\ .$
@@
$H_{2g}(\partial(X_{\Gamma}^{\mathrm{tor}}),\mathbb{Q})=\bigoplus_{{\mathfrak{J}}\in{\mathcal{J}}}H_{2g}({\mathcal{B}}_{\mathfrak{J}},\mathbb{Q})\leavevmode\nobreak\ ,$
@@
$([\widehat{{\mathcal{Z}}}(\underline{\nu},N,Y)].\gamma)_{X^{\mathrm{tor}}_{\Gamma}}=([{\mathcal{Z}}(\underline{\nu},N)_{{\mathfrak{J}}}].\gamma)_{{\mathcal{B}}_{\mathfrak{J}}}-\frac{r_{\mathfrak{J}}}{d_{k}}\sum_{\ell=1}^{g}(\iota_{{\mathfrak{J}},*}[\varphi(\tau,\underline{\nu},N)\cup D^{\ell-1}].\gamma)_{X_{\Gamma}^{\mathrm{tor}}}\leavevmode\nobreak\ ,$
where in the first expression, the intersection is taking place in the boundary divisor ${\mathcal{B}}_{\mathfrak{J}}$. For the second expression, we use Fulton’s excess intersection formula and Corollary 4.5:
$(\iota_{{\mathfrak{J}},*}[\varphi(\tau,\underline{\nu},N)\cup D^{\ell-1}].\gamma)_{X_{\Gamma}^{\mathrm{tor}}}=$ $([\varphi(\tau,\underline{\nu},N)\cup D^{\ell-1}\cup c_{1}({\mathcal{N}}_{{\mathcal{B}}_{\mathfrak{J}}})].\gamma)_{{\mathcal{B}}_{\mathfrak{J}}}$
$=$ $-\frac{d_{k}}{r_{\mathfrak{J}}}([\varphi(\tau,\underline{\nu},N)\cup D^{\ell}].\gamma)_{{\mathcal{B}}_{\mathfrak{J}}}$
By Theorem 4.12, $([\widehat{{\mathcal{Z}}}(\underline{\nu},N,Y)].\gamma)_{X^{\mathrm{tor}}_{\Gamma}}$ is the coefficient of a non-holomorphic Hermitian modular form, which concludes the proof of Theorem 1.3. The proof of Theorem 1.4 on corrected cycles is done similarly using Theorem 4.19.

### v2

@@ 5. Proofs of the Main Theorems
In this section, we prove Theorem˜1.4 and Theorem˜1.3.
Let $(L,h)$ be a Hermitian lattice of signature $(n+1,1)$ over ${\mathcal{O}}_{k}$, the ring of integers of a quadratic imaginary field $k$ and let $X^{\mathrm{tor}}_{\Gamma}$ be the toroidal compactification of the unitary Shimura variety constructed in Section˜3.
Let $g\geq 1$, $N\in\mathrm{Herm}_{g}(k)_{\geq 0}$ and $\underline{\nu}\in(L^{\vee}/L)^{g}$. Let ${\mathcal{Z}}(\underline{\nu},N)$ be the associated special cycle. We again denote its closure in $X^{\mathrm{tor}}_{\Gamma}$ by $\overline{{\mathcal{Z}}}(\underline{\nu},N)$.
For each ${\mathfrak{J}}\subset L$ isotropic rank 11 ${\mathcal{O}}_{k}$-submodule, let $M={\mathfrak{J}}^{\bot}/{\mathfrak{J}}$ and let ${\mathcal{B}}_{\mathfrak{J}}$ be the corresponding boundary divisor of $X^{\mathrm{tor}}_{\Gamma}$. By Lemma˜3.5, the restriction of Kudla–Millson generating series to ${\mathcal{B}}_{\mathcal{J}}$ is the generating series of the $Z(\underline{\lambda})$ cycles introduced in Section˜3.5:
$\Phi_{M}^{g}(\tau)=\sum_{\underline{\lambda}\in(M^{\vee})^{g}}[Z(\underline{\lambda})]q^{h(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\penalty 10000\ .$
We now define the completions of the special cycles: For each ${\mathfrak{J}}\subset L$ isotropic ${\mathcal{O}}_{k}$-line, let $M={\mathfrak{J}}^{\bot}/{\mathfrak{J}}$, which is a Hermitian lattice of signature $(n,0)$. For each $\underline{\nu}\in(M^{\vee}/M)^{g}$ and $N\in\mathrm{Herm}_{g}(k)_{>0}$, we define:
$P_{i}^{\ell}(\underline{\nu},N)=\sum_{\begin{subarray}{c}{\underline{\lambda}\in(M^{\vee})^{g}}\\ [\underline{\lambda}]=\underline{\nu}\\ {(\lambda,\lambda)=N}\end{subarray}}(\Lambda^{g-\ell}P_{i}^{\ell,\ell})(\lambda_{1},\ldots,\lambda_{g})\penalty 10000\ .$
@@
$[\varphi(\tau,\underline{\nu},N)]=\sum_{\underset{[\underline{\lambda}]=\underline{\nu}}{\underline{\lambda}\in(M^{\vee})^{g}}}\sum_{\ell=1}^{g}\frac{(-1)^{\ell}}{(4\pi)^{\ell}(g-\ell)!}\mathrm{Tr}\left(\wedge^{\ell}(Y)\wedge^{g-\ell}([f(\underline{\lambda})]\right)\wedge D_{{\mathfrak{J}}}^{\ell-1}\penalty 10000\ .$
@@
    the completed special cycle in $X_{\Gamma}^{\mathrm{tor}}$: choose a set $\mathcal{J}$} of representative of isotropic lines of $L$ under the action of $\Gamma$ and let
    $\widehat{{\mathcal{Z}}}(\underline{\nu},N,Y)]=[\overline{{\mathcal{Z}}}(\underline{\nu},N)]-\sum_{{\mathfrak{J}}\in{\mathcal{J}}}\frac{r_{\mathfrak{J}}}{d_{k}}\iota_{{\mathfrak{J}},*}[\varphi(\tau,\underline{\nu},N)]\penalty 10000\ .$
@@
    $\widetilde{{\mathcal{Z}}}(\underline{\nu},N)]=[\overline{{\mathcal{Z}}}(\underline{\nu},N)]+\sum_{{\mathfrak{J}}\in{\mathcal{J}}}\frac{r_{\mathfrak{J}}}{d_{k}}\sum_{\underset{i\in I_{\ell}}{\ell=0}}^{g-1}P^{\ell}_{i}(\underline{\nu},N)\,\iota_{{\mathfrak{J}},*}[W^{\ell}_{i}\cup D_{{\mathfrak{J}}}^{g-\ell-1}]\penalty 10000\ ,$
where $d_{k}$ is the discriminant of $k$ and $r_{\mathfrak{J}}$ is the integer appearing in Section˜3.3.
Let $\alpha\in\mathrm{H}_{2g}(X_{\Gamma}^{\mathrm{tor}})$, so we must prove that both pairings $(\widehat{\Phi}_{L}.\alpha)$ and $(\widetilde{\Phi}_{L}.\alpha)$ are Hermitian modular forms of weight $1+n$ and genus $g$.
By Lemma˜3.3, every $\alpha\in\mathrm{H}_{2g}(X_{\Gamma}^{\mathrm{tor}})$ can be written as a sum $\alpha=\beta+\gamma$ where $\beta\in H_{2g}(X_{\Gamma},\mathbb{Q})$ and $\gamma\in\mathrm{H}_{2g}(\partial(X_{\Gamma}^{\mathrm{tor}}),\mathbb{Q})$. In particular, we have:
$\left(\widetilde{\Phi}.\alpha\right)=\left(\widetilde{\Phi}.\beta\right)+\left(\widetilde{\Phi}.\gamma\right)\penalty 10000\ .$
@@
$H_{2g}(\partial(X_{\Gamma}^{\mathrm{tor}}),\mathbb{Q})=\bigoplus_{{\mathfrak{J}}\in{\mathcal{J}}}H_{2g}({\mathcal{B}}_{\mathfrak{J}},\mathbb{Q})\penalty 10000\ ,$
@@
$([\widehat{{\mathcal{Z}}}(\underline{\nu},N,Y)].\gamma)_{X^{\mathrm{tor}}_{\Gamma}}=([{\mathcal{Z}}(\underline{\nu},N)_{{\mathfrak{J}}}].\gamma)_{{\mathcal{B}}_{\mathfrak{J}}}-\frac{r_{\mathfrak{J}}}{d_{k}}\sum_{\ell=1}^{g}(\iota_{{\mathfrak{J}},*}[\varphi(\tau,\underline{\nu},N)\cup D_{{\mathfrak{J}}}^{\ell-1}].\gamma)_{X_{\Gamma}^{\mathrm{tor}}}\penalty 10000\ ,$
where in the first expression, the intersection is taking place in the boundary divisor ${\mathcal{B}}_{\mathfrak{J}}$. For the second expression, we use Fulton’s excess intersection formula and Corollary˜4.5:
$(\iota_{{\mathfrak{J}},*}[\varphi(\tau,\underline{\nu},N)\cup D_{{\mathfrak{J}}}^{\ell-1}].\gamma)_{X_{\Gamma}^{\mathrm{tor}}}=$ $([\varphi(\tau,\underline{\nu},N)\cup D_{{\mathfrak{J}}}^{\ell-1}\cup c_{1}({\mathcal{N}}_{{\mathcal{B}}_{\mathfrak{J}}})].\gamma)_{{\mathcal{B}}_{\mathfrak{J}}}$
$=$ $-\frac{d_{k}}{r_{\mathfrak{J}}}([\varphi(\tau,\underline{\nu},N)\cup D_{{\mathfrak{J}}}^{\ell}].\gamma)_{{\mathcal{B}}_{\mathfrak{J}}}$
By Theorem˜4.12, $([\widehat{{\mathcal{Z}}}(\underline{\nu},N,Y)].\gamma)_{X^{\mathrm{tor}}_{\Gamma}}$ is the coefficient of a non-holomorphic Hermitian modular form, which concludes the proof of Theorem˜1.3. The proof of Theorem˜1.4 on corrected cycles is done similarly using Theorem˜4.19.